OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (dumbcow):

Find factors of (a*c) that add up to b Example: 2x^2 +13x +15 a*c = 2*15 =30 10*3 = 30 10+3 = 13 -> 2x^2 +10x +3x +15 Factor by grouping 2x(x+5) +3(x+5) (2x+3)(x+5)